Light travels through a glass plate of thickness t and refractive index u. If c is the speed of light in vacuum, the time taken by light to travel this thickness of glass is


     


Grade:10

2 Answers

Navjyot Kalra
10 years ago

(d)

        Total thickness = t; Refrative index = m

        Speed of light in Glass plate=c/u

Spandana
4 years ago
A):d
N(refractive index)=c/v
1/n=v/c
V=c/n
Time taken by light to travel the thickness of the glass plate =distance/velocity
                            =t/c/u
                             =ut/c.